# arXiv:2603.01466v1[math.FA] 02 Mar 2026

### v1

@@ 3. Bilocal inequalities and their bounds
Following the assumptions in Section 2, the network correlation $\hat{p}=p(\alpha\beta\gamma|xyz)$ is defined as
@@
where the measurement operators ${A}_{\alpha|x}\in\mathcal{M}_{A},\ {B}_{\beta|y}\in\mathcal{M}_{B},\ {C}_{\gamma|z}\in\mathcal{M}_{C}$, $\{x,y,z\}$ and $\{\alpha,\beta,\gamma\}$ denote the inputs and outputs of measurement operators with $x,y,z,\alpha,\beta,\gamma\in\{0,1\}$. Let $a_{x}\in\mathcal{M}_{A}$, $b_{y}\in\mathcal{M}_{B}$, $c_{z}\in\mathcal{M}_{C}$ be operators such that $I\leq a_{x},b_{y},c_{z}\leq I$ for any $x,y,z\in\{0,1\}$, and $a_{x}=\sum_{\alpha=0}^{1}(-1)^{\alpha}A_{\alpha|x},b_{y}=\sum_{\beta=0}^{1}(-1)^{\beta}B_{\beta|y},c_{z}=\sum_{\gamma=0}^{1}(-1)^{\gamma}C_{\gamma|z}$. Then denote $I_{\tau}=\tau((a_{0}+a_{1})b_{0}(c_{0}+c_{1})),\ J_{\tau}=\tau((a_{0}-a_{1})b_{1}(c_{0}-c_{1}))$. In the MCvNA model, analogous to Ineq. (2.1), we define
@@
Analogous to Ineq. (2.2), under the same assumption, we have the following
(3.2) $\mathcal{S}_{\tau}^{\prime}=|K_{\tau}|+|L_{\tau}|.$
@@
Next, we rely on the abelianness of the algebra to determine the bounds for the bilocal inequality, respectively. We mention that all theorems in this section are based on the independence assumption (*).
The following conclusion indicates that in the mutually-commuting von Neumann algebra model, the maximum upper bound that the quantity (3.1) can achieve is $2\sqrt{2}$. This is analogous to the case in non-relativistic quantum mechanics, where the maximal violation of the bilocal value based on quantum resources is $2\sqrt{2}$.
Theorem 1. In the MCvNA model, $\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}\leq 2\sqrt{2}.$
Proof. According to Gelfand-Namark-Segal (GNS) construction, there is a -representation $\pi_{\tau}:\mathcal{M}_{ABC}\rightarrow H_{\tau}$ and a cyclic vector $\Omega\in H_{\tau}$ such that the set $\{\pi_{\tau}(d)\Omega:\ d\in\mathcal{M}_{ABC}\}$ is dense in $H_{\tau}$. It follows by applying the Cauchy-Schwarz inequality that
$\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}$
@@
$\leq\sqrt{2}\sqrt{\sqrt{\|\pi_{\tau}(b_{0})\Omega\|\|\pi_{\tau}((a_{0}+a_{1})(c_{0}+c_{1}))\Omega\|}+\sqrt{\|\pi_{\tau}(b_{1})\Omega\|\|\pi_{\tau}((a_{0}-a_{1})(c_{0}-c_{1}))\Omega\|}}$
@@
Below, we demonstrate how the value (3.1) depends on the abelianness of the algebras $\mathcal{M}_{A}$, $\mathcal{M}_{B}$ and $\mathcal{M}_{C}$. The results indicate that in entanglement-swapping networks, the abelianness of the three algebras plays distinct roles in setting the upper bound of the inequality to 2. This is not a simple generalization of the bipartite Bell scenario [17], where, with only two systems, Summers et al. showed that if one of these two algebras is abelian, the upper bound of the Bell inequality is 2.
Theorem 2. (1) If $\mathcal{M}_{A}$ and $\mathcal{M}_{C}$ are Abelian, then
$\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}\leq 2.$
(2) If $\mathcal{M}_{B}$ is Abelian, then
$|K_{\tau}|+|L_{\tau}|\leq 2.$
Proof. To see (1), since $\mathcal{M}_{A}$ and $\mathcal{M}_{C}$ are Abelian, respectively, the eight elements
@@
$\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}=\sqrt{|\tau((a_{0}+a_{1})b_{0}(c_{0}+c_{1}))|}+\sqrt{|\tau((a_{0}-a_{1})b_{1}(c_{0}-c_{1}))|}$
@@
where the first inequality follows the fact that $I\leq b_{i}\leq I$ and the order-preserving of state $\tau$, and the second inequality holds because of the Cauchy Schwarz inequality and the non-negativeness of $a_{\epsilon_{0}\epsilon_{1}},\ b_{\epsilon_{0}\epsilon_{1}}$.
To prove (2), since $\mathcal{M}_{B}$ is Abelian, the four elements
@@
Remark. The above theorem illustrates a phenomenon: the violations of the two Bell-type inequalities $\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}\leq 2$ and $|K_{\tau}|+|L_{\tau}|\leq 2$ can indicate the abelianness of the underlying algebras. Conversely, applying the theorem, we can deduce from the violation of an inequality with an upper bound of 2 that the three von Neumann algebras are non-abelian. Here, one also needs to consider the chosen observables and the specific state.
If we wish to set aside the influence of the algebraic structure and consider separately the impact of the choice of observables and the state on values of (3.1) and (3.2), then the following conclusion can be regarded as a generalization of the results in the non-classical mechanical framework. For the sake of clarity, we introduce a tensor product structure into the three-algebra MCvNA model. Let $\mathcal{M}_{B}=\mathcal{M}_{B}^{L}\otimes\mathcal{M}_{B}^{R}$. We call the state $\tau\in\mathcal{M}_{ABC}^{*}$ is separable on $(\mathcal{M}_{A}\vee\mathcal{M}_{B}^{L})^{\prime\prime}$ if there are $\xi_{\alpha}\in\mathcal{M}_{A}^{*},\eta_{\alpha}\in(\mathcal{M}_{B}^{L})^{*}$ such that $\tau(ab^{L})=\sum_{\alpha}\lambda_{\alpha}\xi_{\alpha}(a)\eta_{\alpha}(b^{L})$ with $\sum_{\alpha}\lambda_{\alpha}=1$.
Theorem 3. $\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}\leq 2$ for any state $\tau$ that is separable on $(\mathcal{M}_{A}\vee\mathcal{M}_{B}^{L})^{\prime\prime}$ and $(\mathcal{M}_{B}^{R}\vee\mathcal{M}_{C})^{\prime\prime}$.
Proof. It follows from separability assumption of $\tau$ that
$\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}=\sqrt{|\tau((a_{0}+a_{1})b_{0}(c_{0}+c_{1}))|}+\sqrt{|\tau((a_{0}-a_{1})b_{1}(c_{0}-c_{1}))|}$
@@
$\mathcal{S}(\tau,\mathcal{M}_{A},\mathcal{M}_{B},\mathcal{M}_{C})=\sup_{\{a_{x},b_{y},c_{z}\}}(\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}).$
Combining Theorems 1, 2 and 3, one naturally obtains the following corollary. A similar conclusion can be drawn for the quantity (3.2).
Corollary 4.
(1) $2\leq\mathcal{S}(\tau,\mathcal{M}_{A},\mathcal{M}_{B},\mathcal{M}_{C})\leq 2\sqrt{2}$.
(2) If $\mathcal{M}_{A}$ and $\mathcal{M}_{C}$ are Abelian, then $\mathcal{S}(\tau,\mathcal{M}_{A},\mathcal{M}_{B},\mathcal{M}_{C})=2$.
(3) $\phi\in(\mathcal{M}_{A}\vee\mathcal{M}_{B_{1}}\vee\mathcal{M}_{B_{2}}\vee\mathcal{M}_{C})^{*}$ is convex combination of separable states, then $\mathcal{S}(\phi,\mathcal{M}_{A},\mathcal{M}_{B},\mathcal{M}_{C})=2$.
(4) $|\mathcal{S}(\phi,\mathcal{M}_{A},\mathcal{M}_{B},\mathcal{M}_{C})-\mathcal{S}(\psi,\mathcal{M}_{A},\mathcal{M}_{B},\mathcal{M}_{C})|\leq k\sqrt{\|\phi-\psi\|}$, so, in particular, $\sup_{\{a_{x},b_{y},c_{z}\}}(\sqrt{|I_{\phi}|}+\sqrt{|J_{\phi}|})$ is norm continuous.
Proof. To show (1). Since $\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}=2$ when $a_{0}=a_{1}=b_{0}=c_{0}=c_{1}=I$. And combinig the proof of Theorem 2, we obtain (1).
(2) holds by facts $\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}=2$ when $a_{0}=a_{1}=b_{0}=c_{0}=c_{1}=I$ and Theorem 2.
(3) holds by facts $\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}=2$ when $a_{0}=a_{1}=b_{0}=c_{0}=c_{1}=I$ and Theorem 3.
To prove (4). According to the representations of $I_{\tau}$ and $J_{\tau}$, and the facts that $|\sup x-\sup y|\leq\sup|x-y|$ and triangle inequality, one can obtain that
$\left|\displaystyle\sup_{\{a_{x},b_{y},c_{z}\}}\left(\sqrt{|I_{\phi}|}+\sqrt{|J_{\phi}|}\right)-\displaystyle\sup_{\{a_{x},b_{y},c_{z}\}}\left(\sqrt{|I_{\psi}|}+\sqrt{|J_{\psi}|}\right)\right|$
@@ 4. Maximal violation of bilocal inequalities and algebraic structures
In this section, we aim to point out that the violation of bilocal inequalities, in particular the maximal violation, can reflect the structural properties of the algebra. Here, violation refers to exceeding the upper bound of 2 for the inequality, and maximal violation refers to attaining $2\sqrt{2}$.
@@
Theorem 5. If $\tau$ is faithful, then
@@
if and only if $\tau(a_{i}^{2}a)=\tau(a),\ \tau(b_{i}^{2}b)=\tau(b),\ \tau(c_{i}^{2}c)=\tau(c)$, and $\tau[(a_{0}a_{1}+a_{1}a_{0})a]=0,\ \tau[(c_{0}c_{1}+c_{1}c_{0})c]=0$ for any $a\in\mathcal{M}_{A},b\in\mathcal{M}_{B},c\in\mathcal{M}_{C}$.
Remark. This theorem states that when the inequality is maximally violated, $I,\ a_{0},\ a_{1}$ and $\frac{i}{2}[a_{0},a_{1}]$ form a realization of the Pauli spin matrices in $\mathcal{M}_{A}$ and generate a copy of $M_{2}(\mathbb{C})$ in $\mathcal{M}_{A}$. And the same holds for $I,\ c_{0},\ c_{1}$ and $\frac{i}{2}[c_{0},c_{1}]$ in $\mathcal{M}_{C}$. Moreover, $\tau(a_{0}a_{1}+a_{1}a_{0})=0,\ \tau(c_{0}c_{1}+c_{1}c_{0})=0.$
@@
From the (4.7), one gets $a_{i}^{2}=I,\ c_{i}^{2}=I\ (i=0,1)$ because $I\leq a_{i},c_{i}\leq I$, therefore $\tau(a_{i}^{2}a)=\tau(a)\ \tau(c_{i}^{2}c)=\tau(c)$ for any $a\in\mathcal{M}_{A},\ b\in\mathcal{M}_{B},\ c\in\mathcal{M}_{C}$.
Now let us show the proof for $\tau(b_{i}^{2}b)=\tau(b),\ \tau[(a_{0}a_{1}+a_{1}a_{0})a]=0$, and $\tau[(c_{0}c_{1}+c_{1}c_{0})c]=0$ for any $a\in\mathcal{M}_{A},\ b\in\mathcal{M}_{B},\ c\in\mathcal{M}_{C}$, which implies $b_{i}^{2}=I,\ a_{0}a_{1}+a_{1}a_{0}=0,$ and $c_{0}c_{1}+c_{1}c_{0}=0$. According to the Eqs. (4.5)-(4.7), i.e.,
@@
Then combing the conditions (4.2), (4.3) and condition (4.4) one can get
@@
deriving that $\tau(X)=\tau(Y)=0$. It follows from $\tau(X)=0=\tau(Y)$ and $\|\pi_{\tau}(b_{i})\Omega\|=1$ that
@@
Similarly $\tau(b_{1}^{2}b)=\tau(b)$ for any $b\in\mathcal{M}_{B}.$ Furthermore, it follows from Eq. (4.2) and $\pi_{\tau}[(b_{0}^{2})]\Omega=\Omega$ that
@@
The above Theorem 5 indicates that if the bilocal inequality is maximally violated, then the underlying algebra locally contains an $M_{2}(\mathbb{C})$ factor. However, this does not lead to further global structural properties of the algebra. Below, we aim to further explore the relationship between the maximal violation of the inequality and the global properties of the algebra.
@@
Theorem 6. Let $\mathcal{M}_{ABC}$ be a von Neumann algebra on a separable Hilbert space $H$ with cyclic and separating vector. If there are hyperfinite type II1 factors $\mathcal{R}_{A}$ and $\mathcal{R}_{C}$ such that $\mathcal{M}_{A}\simeq\mathcal{M}_{A}\otimes\mathcal{R}_{A}$ and $\mathcal{M}_{C}\simeq\mathcal{M}_{C}\otimes\mathcal{R}_{C}$, then $\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}=2\sqrt{2}$ for every normal state $\tau$.

### v2

@@ 4. Bilocal inequalities and their bounds
In this section, we further analyze the conditions under which the bilocal inequality holds or is violated in the TMCvNA model of an entanglement swapping network. Specifically, in this model, we can construct bilocal inequalities analogous to those in the non-relativistic setting. Define
$I_{\tau}=\tau\bigl((A_{0}+A_{1})B_{0}(C_{0}+C_{1})\bigr),\ J_{\tau}=\tau\bigl((A_{0}-A_{1})B_{1}(C_{0}-C_{1})\bigr),$
where $\tau$ is the state on $\mathcal{M}_{\mathcal{ABC}}$ satisfying the independent condition (*). Here, $\tau\bigl(A_{x}B_{y}C_{z}\bigr)=\sum_{a,b,c=0}^{1}(-1)^{a+b+c}\tau({A}_{a|x}{B}_{b|y}{C}_{c|z})$ and $A_{x}=\sum_{a}(-1)^{a}A_{a|x}$, $B_{y}=\sum_{b}(-1)^{b}B_{b|y}$, $C_{z}=\sum_{z}(-1)^{z}C_{c|z}$, where $A_{a|x}$, $B_{b|y}$, and $C_{c|z}$ are the POVMs performed by Alice, Bob, and Charles, respectively. Moreover, the network correlation $\hat{p}=p(\alpha\beta\gamma|xyz)$ in the TMCvNA model is defined as
@@
In the TMCvNA model, analogous to Ineq. (3.1), we set
@@
Analogously, corresponding to Ineq. (3.2), we define
(4.2) $\mathcal{S}_{\tau}^{\prime}=|K_{\tau}|+|L_{\tau}|,$
@@
We say that the state $\tau$ together with the observables $A_{x}$, $B_{y}$, $C_{z}$ satisfies the bilocal inequality if $\mathcal{S}_{\tau}\leq 2$, and violates it if $2<\mathcal{S}_{\tau}\leq 2\sqrt{2}$. The same criterion applies to $\mathcal{S}_{\tau}^{\prime}$ in Ineq. (4.2).
Next, we rely on the abelianness of the algebra to determine the bounds for the bilocal inequality, respectively.
The following conclusion indicates that in the mutually-commuting von Neumann algebra model, the supremum of $\mathcal{S}_{\tau}$ defined in Ineq. (4.1) is $2\sqrt{2}$. This coincides with the case in non-relativistic quantum mechanics, where the bilocal quantity $\mathcal{S}$ in Ineq. (3.1) attains a maximal violation of $2\sqrt{2}$ based on quantum resources.
Theorem 1. In the ternary mutually-commuting von Neumann algebra models of entanglement swapping networks, $\mathcal{S}_{\tau}=\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}\leq 2\sqrt{2}.$
Proof. According to the Gelfand-Namark-Segal (GNS) construction, there is a -representation $\pi_{\tau}:\mathcal{M}_{\mathcal{ABC}}\rightarrow\mathcal{B}(\mathcal{H}_{\tau})$ and a cyclic vector $\Omega\in\mathcal{B}(\mathcal{H}_{\tau})$ such that the set $\{\pi_{\tau}(O)\Omega:\ O\in\mathcal{M}_{\mathcal{ABC}}\}$ is dense in $\mathcal{H}_{\tau}$. It follows by applying the Cauchy-Schwarz inequality that
$\mathcal{S}_{\tau}=\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}$
@@
$\leq\sqrt{2}\sqrt{\|\pi_{\tau}(B_{0})\Omega\|\|\pi_{\tau}((A_{0}+A_{1})(C_{0}+C_{1}))\Omega\|+\|\pi_{\tau}(B_{1})\Omega\|\|\pi_{\tau}((A_{0}-A_{1})(C_{0}-C_{1}))\Omega\|}$
@@
Building on the results above, we now investigate how the quantity $\mathcal{S}_{\tau}$ in Ineq. (4.1) depends on the abelianness of the algebras $\mathcal{M}_{\mathcal{A}}$, $\mathcal{M}_{\mathcal{B}}$, and $\mathcal{M}_{\mathcal{C}}$. Specifically, the results indicate that in entanglement swapping networks, the abelianness of the three algebras plays distinct roles in reducing the upper bound of the inequality to 2, i.e., determining the conditions under which no violation of the bilocal inequality can occur. This is not a simple generalization of the bipartite Bell scenario [27], where, with only two systems, Summers et al. showed that if one of these two algebras is abelian, the upper bound of the Bell inequality is 2.
Theorem 2. In the ternary mutually-commuting von Neumann algebra models of entanglement swapping networks,
(1) if $\mathcal{M}_{\mathcal{A}}$ and $\mathcal{M}_{\mathcal{C}}$ are Abelian, then
$\mathcal{S}_{\tau}=\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}\leq 2;$
(2) if $\mathcal{M}_{\mathcal{B}}$ is Abelian, then
$\mathcal{S}_{\tau}^{\prime}=|K_{\tau}|+|L_{\tau}|\leq 2.$
Proof. To see (1), since $\mathcal{M}_{\mathcal{A}}$ and $\mathcal{M}_{\mathcal{C}}$ are Abelian, respectively, the eight elements
@@
$\mathcal{S}_{\tau}=\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}=\sqrt{|\tau((A_{0}+A_{1})B_{0}(C_{0}+C_{1}))|}+\sqrt{|\tau((A_{0}-A_{1})B_{1}(C_{0}-C_{1}))|}$
@@
where the first inequality follows the fact that $I\leq B_{i}\leq I\ (i=0,1)$ and the order-preserving of state $\tau$, and the second inequality holds because of the Cauchy Schwarz inequality and the non-negativeness of $A_{\epsilon_{0}\epsilon_{1}}\ \text{and}\ \ B_{\epsilon_{0}\epsilon_{1}}$.
To prove (2), since $\mathcal{M}_{\mathcal{B}}$ is Abelian, the four elements
@@
Remark. The above theorem illustrates a phenomenon: the violations of the two Bell-like inequalities, i.e., $2<\mathcal{S}_{\tau}\leq 2\sqrt{2}$ and $2<\mathcal{S}^{\prime}_{\tau}\leq 2\sqrt{2}$ can serve as an indicator of the non-abelianness of the underlying algebras. Applying the theorem, we can infer from a violation of Ineq. (4.1) that at least one of the algebras $\mathcal{M_{A}}$ and $\mathcal{M_{C}}$ is non-abelian; Similarly, a violation of Ineq. (4.2) implies that $\mathcal{M_{B}}$ is non-abelian. Here, one also needs to consider the chosen observables and the specific state.
Nevertheless, it is important to note that these conclusions depend not only on the algebraic structure but also on the specific choice of observables and the particular state under consideration. To isolate the influence of the algebraic structure and examine separately the role of observables and the state in determining the values $\mathcal{S}_{\tau}$ and $\mathcal{S}_{\tau}^{\prime}$ in Ineqs. (4.1) and  (4.2), the following conclusion can be regarded as a generalization of the results in the non-relativistic quantum mechanical framework. For the sake of clarity, we introduce a tensor product structure into the TMCvNA model. Before presenting the following result, we first introduce the definition of separable states.
Definition 3. (Separable states) Let $\mathcal{M}_{\mathcal{A}}$ and $\mathcal{M}_{\mathcal{B}}$ be mutually commuting von Neumann subalgebras of $\mathcal{M}_{\mathcal{AB}}=(\mathcal{M}_{\mathcal{A}}\vee\mathcal{M}_{\mathcal{B}})^{\prime\prime}$. A state $\tau\in\mathcal{M}_{\mathcal{AB}}^{*}$ is said to be separable if there exist states $\xi_{\alpha}\in\mathcal{M}_{\mathcal{A}}^{*}$, $\eta_{\alpha}\in\mathcal{M}_{\mathcal{B}}^{*}$, and positive coefficients $\lambda_{\alpha}$ with $\sum_{\alpha}\lambda_{\alpha}=1$ such that for all $A\in\mathcal{M}_{\mathcal{A}},B\in\mathcal{M}_{\mathcal{B}}$,
$\tau(AB)=\sum_{\alpha}\lambda_{\alpha}\xi_{\alpha}(A)\eta_{\alpha}(B).$
[…]
$\sup_{\{A_{x},B_{y},C_{z}\}}\mathcal{S}_{\tau}=\displaystyle\sup_{\{A_{x},B_{y},C_{z}\}}(\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|})=2.$
[…]
$\mathcal{S}_{\tau}=\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}=\sqrt{|\tau((A_{0}+A_{1})B_{0}(C_{0}+C_{1}))|}+\sqrt{|\tau((A_{0}-A_{1})B_{1}(C_{0}-C_{1}))|}$
@@
$\mathcal{S}(\tau,\mathcal{M}_{\mathcal{A}},\mathcal{M}_{\mathcal{B}},\mathcal{M}_{\mathcal{C}})=\sup_{\{A_{x},B_{y},C_{z}\}}(\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|})$
that appears in Theorem 3.
Combining Theorems 1, 2 and 3, one naturally obtains the following corollary. A similar conclusion can be drawn for the quantity $\mathcal{S}^{\prime}_{\tau}$ in Ineq. (4.2).
Corollary 4. In the ternary mutually-commuting von Neumann algebra models of entanglement swapping networks,
(1) for any state $\tau$ and any choice of observables $A_{x}\in\mathcal{M}_{\mathcal{A}},\ B_{y}\in\mathcal{M}_{\mathcal{B}},\ C_{z}\in\mathcal{M}_{\mathcal{C}}$ in a scheme with two inputs and two outputs, the quantity $\mathcal{S}(\tau,\mathcal{M}_{A},\mathcal{M}_{B},\mathcal{M}_{C})$ satisfies
$2\leq\mathcal{S}(\tau,\mathcal{M}_{\mathcal{A}},\mathcal{M}_{\mathcal{B}},\mathcal{M}_{\mathcal{C}})\leq 2\sqrt{2}.$
(2) if $\mathcal{M}_{\mathcal{A}}$ and $\mathcal{M}_{\mathcal{C}}$ are Abelian, then $\mathcal{S}(\tau,\mathcal{M}_{\mathcal{A}},\mathcal{M}_{\mathcal{B}},\mathcal{M}_{\mathcal{C}})=2$.
(3) for any states $\phi,\psi\in[(\mathcal{M}_{\mathcal{A}}\vee\mathcal{M}_{\mathcal{B}}\vee\mathcal{M}_{\mathcal{C}})^{\prime\prime}]^{*}$, the following inequality holds:
$|\mathcal{S}(\phi,\mathcal{M}_{\mathcal{A}},\mathcal{M}_{\mathcal{B}},\mathcal{M}_{\mathcal{C}})-\mathcal{S}(\psi,\mathcal{M}_{\mathcal{A}},\mathcal{M}_{\mathcal{B}},\mathcal{M}_{\mathcal{C}})|\leq k\sqrt{\|\phi-\psi\|},$
where $k$ is a positive constant. Consequently, the functional $\phi\rightarrow\mathcal{S}(\phi,\mathcal{M}_{\mathcal{A}},\mathcal{M}_{\mathcal{B}},\mathcal{M}_{\mathcal{C}})=\displaystyle\sup_{\{A_{x},B_{y},C_{z}\}}(\sqrt{|I_{\phi}|}+\sqrt{|J_{\phi}|})$ is norm continuous.
Proof. To show (1), note that $\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}=2$ when $A_{0}=A_{1}=B_{0}=C_{0}=C_{1}=I$, and combining this with the proof of Theorem 2, we obtain (1).
(2) holds by the fact that $\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}=2$ when $A_{0}=A_{1}=B_{0}=C_{0}=C_{1}=I$, and by Theorem 2.
To prove (3), note that by the representations of $I_{\tau}$ and $J_{\tau}$, together with the facts that $|\sup x-\sup y|\leq\sup|x-y|$ and the triangle inequality, one can obtain
$|\mathcal{S}(\phi,\mathcal{M}_{\mathcal{A}},\mathcal{M}_{\mathcal{B}},\mathcal{M}_{\mathcal{C}})-\mathcal{S}(\psi,\mathcal{M}_{\mathcal{A}},\mathcal{M}_{\mathcal{B}},\mathcal{M}_{\mathcal{C}})|$
$=\left|\displaystyle\sup_{\{A_{x},B_{y},C_{z}\}}\left(\sqrt{|I_{\phi}|}+\sqrt{|J_{\phi}|}\right)-\displaystyle\sup_{\{A_{x},B_{y},C_{z}\}}\left(\sqrt{|I_{\psi}|}+\sqrt{|J_{\psi}|}\right)\right|$
@@ 5. Maximal violation of bilocal inequalities and algebraic structures
In this section, we aim to point out that the violation of bilocal inequalities, in particular the maximal violation, can reflect the structural properties of the algebra. Here, violation refers to exceeding the upper bound of 2 for the $\mathcal{S}_{\tau}$ in inequality (4.1), while maximal violation means attaining the value $2\sqrt{2}$ for the same quantity.
@@
Theorem 5. In the ternary mutually-commuting von Neumann algebra models of entanglement swapping networks, if $\tau$ is faithful, then
@@
if and only if $\tau(A_{i}^{2}A)=\tau(A),\ \tau(B_{i}^{2}B)=\tau(B),\ \tau(C_{i}^{2}C)=\tau(C)$, and $\tau[(A_{0}A_{1}+A_{1}A_{0})A]=0,\ \tau[(C_{0}C_{1}+C_{1}C_{0})C]=0$ for any $A\in\mathcal{M}_{\mathcal{A}},\ B\in\mathcal{M}_{\mathcal{B}},\ C\in\mathcal{M}_{\mathcal{C}}$ with $i\in\{0,1\}$.
Remark. This theorem states that when the inequality is maximally violated, $A_{0},\ A_{1}$ and $\frac{i}{2}[A_{0},A_{1}]$ form a realization of the Pauli spin matrices in $\mathcal{M}_{\mathcal{A}}$ and generate a copy of $M_{2}(\mathbb{C})$ in $\mathcal{M}_{\mathcal{A}}$. The same holds for $C_{0},\ C_{1}$ and $\frac{i}{2}[C_{0},C_{1}]$ in $\mathcal{M}_{\mathcal{C}}$. Moreover, $\tau(A_{0}A_{1}+A_{1}A_{0})=0,\ \tau(C_{0}C_{1}+C_{1}C_{0})=0.$
@@
From (5.7), one gets $A_{i}^{2}=I,\ C_{i}^{2}=I\ (i=0,1)$ because $I\leq A_{i},C_{i}\leq I$, therefore $\tau(A_{i}^{2}A)=\tau(A),\ \tau(C_{i}^{2}C)=\tau(C)$ for any $A\in\mathcal{M}_{\mathcal{A}},\ B\in\mathcal{M}_{\mathcal{B}},\ C\in\mathcal{M}_{\mathcal{C}}$.
Now, let us show the proof for $\tau(B_{i}^{2}B)=\tau(B),\ \tau[(A_{0}A_{1}+A_{1}A_{0})A]=0$, and $\tau[(C_{0}C_{1}+C_{1}C_{0})C]=0$ for any $A\in\mathcal{M}_{\mathcal{A}},\ B\in\mathcal{M}_{\mathcal{B}},\ C\in\mathcal{M}_{\mathcal{C}}$, which implies $B_{i}^{2}=I,$ $A_{0}A_{1}+A_{1}A_{0}=0,$ and $C_{0}C_{1}+C_{1}C_{0}=0$. According to Eqs. (5.5)-(5.7), i.e.,
@@
Then combing conditions (5.2), (5.3) and condition (5.4), one can get
@@
deriving $\tau(X)=\tau(Y)=0$. It follows from $\tau(X)=0=\tau(Y)$ and $\|\pi_{\tau}(B_{i})\Omega\|=1$ that
@@
Similarly $\tau(B_{1}^{2}B)=\tau(B)$ for any $B\in\mathcal{M}_{\mathcal{B}}.$ Furthermore, it follows from Eq. (5.2) and $\pi_{\tau}[(B_{0}^{2})]\Omega=\Omega$ that
@@
The above Theorem 5 indicates that if the bilocal inequality is maximally violated, then the underlying algebra locally contains an $M_{2}(\mathbb{C})$ factor. In contrast, the follwing theorem presents a highly specialized scenario in which a precise characterization of algebras attaining maximal viloation can be obtained. Note that next we denote by $\Lambda_{1}\overline{\otimes}\Lambda_{2}$ the von Neumann algebra of the tensor product of $\Lambda_{1},\Lambda_{2}$
@@
Theorem 6. In the ternary mutually-commuting von Neumann algebra models of entanglement swapping networks, if there are hyperfinite type II1 factors $\mathcal{R}_{\mathcal{A}}$ and $\mathcal{R}_{\mathcal{C}}$ such that $\mathcal{M}_{\mathcal{A}}\simeq\mathcal{M}_{\mathcal{A}}\overline{\otimes}\mathcal{R}_{\mathcal{A}}$ and $\mathcal{M}_{\mathcal{C}}\simeq\mathcal{M}_{\mathcal{C}}\overline{\otimes}\mathcal{R}_{\mathcal{C}}$, then $\sqrt{|I_{\tau}|}+\sqrt{|J_{\tau}|}=2\sqrt{2}$ for every normal state $\tau$.